if the perimeter of a rectangle backyard is 6x+6 and the width is x yards find the binomial that represents the length
Answer by stanbon(75887)

if the perimeter of a rectangle backyard is 6x+6 and the width is x yards find the binomial that represents the length
-----------------------
Perimeter = 2(L+W)
6x+6y = 2(L + x)
6x + 6= 2L + 2x
2L = 4x + 6
Length = 2x + 3
======================
Cheers,
Stan H.